OpenStudy (anonymous):

he number of men in a tour group is 15 times the number of women. If there are 48 people in the group, how many are men? 45 30 15 10

OpenStudy (k12awesomeness):

I can help you

OpenStudy (k12awesomeness):

The answer is 45... because,System: 1) x+y=48 2) x=15y Solution... (15y)+y=48 y=3 x=15*3=45 45 men

OpenStudy (k12awesomeness):

Does this help?

OpenStudy (k12awesomeness):

Well, I hope it does help... Tag me if you need me